a sphere of mass 0.5kg and diameter 1m rolls without slipping with a constant velocity of 5ms-1. calculate its total energy
1
Expert's answer
2020-10-12T07:48:16-0400

Express its kinetic energy of linear motion of its center of mass:


"E_L=\frac{mv^2}{2}."

Express its angular velocity in terms of linear velocity and radius:


"\omega=v/r."


Express the moment of inertia:


"I=\frac{2}{5}mr^2."

Express the kinetic energy of rotational motion:


"E_R=\frac{I\omega^2}{2.}"

The total energy of motion is


"E=E_L+E_R,\\\space\\
E=\frac{mv^2}{2}+\frac{mr^2v^2}{5r^2}=\frac{7}{10}mv^2=8.75\text{ J}."